### v1

@@ 4.3 Properties of the construction
Observe that for any $\tau$ that satisfies Proposition 4.6, it holds that every cell of $B+\tau$ lies in between two consecutive cells of $R$. Note that the consecutive red cells are at the shorter distance than consecutive blue cells. Hence, in between two consecutive red cells lies at most one blue cell.
Next, observe that each red cell consists of $\geq 16d$ points. Each blue cell consists of just $2d$ points. Therefore, each point cell of $B+\tau$ is matched to a point in either the next left or the next right red cell. The distance between the $\ell^{\mathrm{th}}$ red cell and the $k^{\mathrm{th}}$ blue cell is:
@@
Observe that when $\lvert k-j_{\tau}\rvert<n/2$ this is minimized for $\ell=2n+i_{\tau}-j_{\tau}+k$. When $\lvert k-j_{\tau}\rvert>n/2$ we have two cases based on the sign of $(k-j_{\tau})$ which matches the definition of the $\textup{NearestCell}(k)$ cell because $|\varepsilon|<0.1$. ∎
@@ 4.4 Equivalence
We assume that $\tau$ is of form defined in Proposition 4.6 as otherwise $\mathrm{EMD}(B+\tau,R)$ is large. Moreover, by 4.8 we can assume that $|\varepsilon|\leq 0.1$. The distance between the left endpoint of the $k^{\mathrm{th}}$ cell of $B$ and the leftmost endpoint of the nearest cell of $R$ is given by:
@@
Because each pair of vectors in the instance is not orthogonal, by Lemma 4.5, we know that the smallest distance between points in each cell is at least $1$. Hence, after plugging in the definition of $\textup{NearestCell}(k)$ and by Lemma 4.5, the contribution of points in the $k^{\mathrm{th}}$ cell of $B$ to $\mathrm{EMD}(B+\tau,R)$ is:
$\mathrm{Cost}(k)\geq\begin{cases}1&\text{if }k=j_{\tau},\\ |j_{\tau}-k+\varepsilon|\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j_{\tau}\rvert<n/2,\\ (n-1-|j_{\tau}-k+\varepsilon|)\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j_{\tau}\rvert>n/2.\\ \end{cases}$
Observe that $\mathrm{Cost}(k)$ is at least $c_{1}\Delta\cdot\min\{\lvert j_{\tau}-k+\varepsilon\rvert,\lvert n-1-j_{\tau}+k-\varepsilon\rvert,\lvert n-1+j_{\tau}-k+\varepsilon\rvert\}+c_{2}$ when $k\neq j_{\tau}$. Hence, the total length of $\mathrm{EMD}(B+\tau,R)$ is at least:
$\sum_{k=0}^{n-1}\mathrm{Cost}(k)\geq$ $1+(n-1)c_{2}+$
$c_{1}\Delta\left(\sum_{k\in\{0,\ldots,n-1\}\setminus\{j_{\tau}\}}\min\{\lvert j_{\tau}-k+\varepsilon\rvert,\lvert n-1-j_{\tau}+k-\varepsilon\rvert,\lvert n-1+j_{\tau}-k+\varepsilon\rvert\}\right).$
Next, let us focus on the last term. We use Proposition 4.7 to conclude that:
$\sum_{k=0}^{n-1}\mathrm{Cost}(k)\geq 1+c_{2}(n-1)+c_{1}\Delta\cdot(n^{2}-1)/4=1+\Lambda.\qed$
@@
Take $\tau^{\ast}=((n-1)(2n+i)-nj)\cdot\Delta$. Observe that the $(2n+i)^{\mathrm{th}}$ cell of $R$ and the $j^{\mathrm{th}}$ cell of $B$ align. Because $x_{i}$ and $y_{j}$ are orthogonal, Lemma 4.5 guarantees that $\mathrm{EMD}(R(x_{i}),B(y_{j}))=0$. By 4.9, all remaining cells are matched to their closest cell. Therefore, by Lemma 4.5 the contribution of edges with endpoints in the $k^{\mathrm{th}}$ cell of $B$ to $\mathrm{EMD}(B+\tau^{\ast},R)$ is
@@
for any $k\in\{0,\ldots,n-1\}\setminus\{j\}$. By plugging in the exact values of $\textup{NearestCell}(k)$, with the similar arguments as in the previous proof, we conclude that:
$\mathrm{Cost}(k)$ =$=\begin{cases}0&\text{if }k=j,\\ \lvert j-k\rvert\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j\rvert<n/2,\\ (n-1-\lvert j-k\rvert)\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j\rvert\geq n/2.\end{cases}$
@@
$\sum_{k=0}^{n-1}\mathrm{Cost}(k)$ $=\sum_{k\in\{0,\ldots,n-1\}\setminus\{j\}}c_{2}+\Delta c_{1}\cdot\min\{\lvert j-k\rvert,\lvert n-1-j+k\rvert,\lvert n-1+j-k\rvert\}$
$=c_{2}(n-1)+2\Delta c_{1}\cdot(n^{2}-1)/4=\Lambda.$
where in the last inequality we used Proposition 4.7. ∎
@@ 5 Lower bounds in higher dimension
In this section, we prove conditional lower bounds for approximating $\mathrm{EMDuT}$ with the $L_{1}$ or $L_{\infty}$ norm. Our lower bounds assume the popular Exponential Time Hypothesis (ETH), which postulates that the 3-SAT problem on $N$ variables cannot be solved in time $2^{o(N)}$ [29].
@@
Now we provide the proof details. Let $\cal{B}$ be the bounding box of $\bigcup_{i=1}^{k}B_{i}\cup R_{i}$, and let $\Delta$ be the sum of all side lengths of $\cal{B}$ (i.e., $\Delta$ is the $L_{1}$ diameter of $\cal{B}$). We set $U\coloneqq(2n+5)\Delta$. We construct the sets $B,R$ as
@@ 5.1 Lower bound for $L_{1}$ asymmetric
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})\geq 2\min_{(u,v)\in E}\Big{(}|\tau_{i}-u|+|\tau_{j}-v|\Big{)}+(d-2)N\geq(d-2)N+1.\qed$
We apply the Gadget Combination Lemma to the gadgets $B_{i,j},R_{i,j},B^{\prime}_{i,j},R^{\prime}_{i,j}$ for $1\leq i<j\leq d$. The $\mathrm{EMDuT}_{1}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the above lemma implies the following. If $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k})\in[N]^{d}$ has a total cost of (${d\choose 2}\cdot(d-2)N\eqcolon\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Thus, each pair of gadgets contributes cost at least $(d-2)N$, and at least one pair of gadgets contributes cost at least $(d-2)N+1$, so the total cost is at least (${d\choose 2}\cdot(d-2)N+1=\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we would have a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2})$, $1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N)$, and $d=k$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{3})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.
@@ 5.2 Lower Bound for $L_{1}$ Symmetric
$(\bar{p}_{i,u,j,v,b})_{\ell}=\begin{cases}u&\text{if }\ell\in\{i,i+k\},\\ v&\text{if }\ell\in\{j,j+k\},\\ b&\text{otherwise,}\end{cases}\qquad(q_{i,j})_{\ell}=\begin{cases}N&\text{if }\ell\in\{i,j\},\\ -N&\text{if }\ell\in\{i+k,j+k\},\\ 0&\text{otherwise.}\end{cases}$
We construct the following $2{k\choose 2}$ gadgets. For any $1\leq i<j\leq k$ we construct [^4] Here, we treat $B_{i,j}$ as a multi-set, containing $|E|-1$ times the points $q_{i,j}$. This can be avoided by adding a tiny perturbation to each copy, which makes $B_{i,j}$ a set without significantly changing any distances.
@@
Note that in any dimension $\ell\not\in\{i,j,i+k,j+k\}$ all points in $B_{i,j}$ and $B^{\prime}_{i,j}$ have coordinate 0, and thus the contribution of these dimensions to the total cost does not depend on the matching. For $(B_{i,j},R_{i,j})$ such a dimension $\ell$ contributes $|E|\cdot|\tau_{\ell}|\geq|E|\tau_{\ell}$ to the total cost, and for $(B^{\prime}_{i,j},R^{\prime}_{i,j})$ it contributes $|E|\cdot|N-\tau_{\ell}|\geq|E|(N-\tau_{\ell})$. Summing up over both gadgets and all $\ell\not\in\{i,j,i+k,j+k\}$ yields total cost at least $(d-4)|E|N$, with equality if $\tau\in[N]^{d}$.
@@
where we used $|x|\geq x$. Furthermore, equality holds if $\tau\in[N]^{d}$. Analogously, one can show that projected onto the dimensions $i,j,i+k,j+k$ the $L_{1}$ distance from $-q_{i,j}+\tau$ to $\bar{p}_{i,u,j,v,N}$ is
@@
By summing up everything, and noting that the terms $\tau_{i}+\tau_{j}-\tau_{i+k}-\tau_{j+k}$ and $-\tau_{i}-\tau_{j}+\tau_{i+k}+\tau_{j+k}$ cancel, we obtain
@@
We apply the Gadget Combination Lemma to the gadgets $B_{i,j},R_{i,j},B^{\prime}_{i,j},R^{\prime}_{i,j}$ for $1\leq i<j\leq k$. The $\mathrm{EMDuT}_{1}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the above lemma implies the following. If $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k},v_{1},\ldots,v_{k})\in[N]^{d}$ has a total cost of (${k\choose 2}\cdot((d+4)|E|-8)N=:\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Thus, each pair of gadgets contributes cost at least $((d+4)|E|-8)N$, and at least one pair of gadgets contributes cost at least $((d+4)|E|-8)N+1$, so the total cost is at least (${k\choose 2}\cdot((d+4)|E|-8)N+1=\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we would have a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2}),1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N^{3})$, and $d=\mathcal{O}(k)$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{5})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.

### v2

@@ 4.3 Properties of the Construction
Observe that for any $\tau$ that satisfies Proposition˜4.6, it holds that every cell of $B+\tau$ lies between two consecutive cells of $R$. Note that the consecutive red cells are at a shorter distance than the consecutive blue cells. Hence, in between two consecutive red cells lies at most one blue cell.
Next, observe that each red cell consists of $\geq 16d$ points. Each blue cell consists of just $2(d+1)$ points. Therefore, each point cell of $B+\tau$ is matched to a point in either the next left or the next right red cell. The distance between the $\ell^{\mathrm{th}}$ red cell and the $k^{\mathrm{th}}$ blue cell is:
@@
Observe that when $\lvert k-j_{\tau}\rvert<n/2$ this is minimized for $\ell=2n+i_{\tau}-j_{\tau}+k$ because $|\varepsilon|<0.1$. When $\lvert k-j_{\tau}\rvert\geq n/2$ we have two cases based on the sign of $(k-j_{\tau})$ which matches the definition of the $\textup{NearestCell}(k)$ cell because $|\varepsilon|<0.1$. ∎
@@ 4.4 Equivalence
We assume that $\tau$ is of form defined in Proposition˜4.6 as otherwise $\mathrm{EMD}(B+\tau,R)$ is large. Moreover, by ˜4.8 we can assume that $|\varepsilon|\leq 0.1$. The distance between the left endpoint of the $k^{\mathrm{th}}$ cell of $B$ and the leftmost endpoint of the nearest cell of $R$ is given by:
@@
Because each pair of vectors in the instance is not orthogonal, by Lemma˜4.5, we know that the smallest distance between points in each cell is at least $1$. Hence, after plugging in the definition of $\textup{NearestCell}(k)$ and by Lemma˜4.5, the contribution of points in the $k^{\mathrm{th}}$ cell of $B$ to $\mathrm{EMD}(B+\tau,R)$ is:
[…]
Observe that $\mathrm{Cost}(k)$ is at least $c_{1}\Delta\cdot\min\{\lvert j_{\tau}-k+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert,\lvert j_{\tau}-k+(n-1)+\varepsilon\rvert\}+c_{2}$ when $k\neq j_{\tau}$. Hence, the total length of $\mathrm{EMD}(B+\tau,R)$ is at least $\sum_{k=1}^{n-1}\mathrm{Cost}(k)$ which is bounded by:
$1+(n-2)c_{2}+c_{1}\Delta\Big(\sum_{k\in\{1,\ldots,n-1\}\setminus\{j_{\tau}\}}\min\{$ $\lvert j_{\tau}-k+\varepsilon\rvert,\lvert k-j_{\tau}-(n-1)+\varepsilon\rvert$,
[…]
We use Proposition˜4.7 to conclude that:
$\sum_{k=1}^{n-1}\mathrm{Cost}(k)\geq 1+c_{2}(n-2)+c_{1}\Delta\cdot n(n-2)/4=1+\Lambda.\qed$
@@
Take $\tau^{\ast}=((n-1)(2n+i)-nj)\cdot\Delta$. Observe that the $(2n+i)^{\mathrm{th}}$ cell of $R$ and the $j^{\mathrm{th}}$ cell of $B$ align. Because $x_{i}$ and $y_{j}$ are orthogonal, Lemma˜4.5 guarantees that $\mathrm{EMD}(R(x_{i}),B(y_{j}))=0$. By ˜4.9, all remaining cells are matched to their closest cell. Therefore, by Lemma˜4.5 the contribution of edges with endpoints in the $k^{\mathrm{th}}$ cell of $B$ to $\mathrm{EMD}(B+\tau^{\ast},R)$ is
@@
for any $k\in\{1,\ldots,n-1\}\setminus\{j\}$. By plugging in the exact values of $\textup{NearestCell}(k)$, with similar arguments as in the previous proof, we conclude that:
$\mathrm{Cost}(k)$ $=\begin{cases}0&\text{if }k=j,\\ \lvert j-k\rvert\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j\rvert<n/2,\\ (n-1-\lvert j-k\rvert)\cdot c_{1}\Delta+c_{2}&\text{if }\lvert k-j\rvert\geq n/2.\end{cases}$
@@
$\sum_{k=1}^{n-1}\mathrm{Cost}(k)$ $=\sum_{k\in\{1,\ldots,n-1\}\setminus\{j\}}c_{2}+\Delta c_{1}\cdot\min\{\lvert j-k\rvert,\lvert n-1-j+k\rvert,\lvert n-1+j-k\rvert\}$
$=c_{2}(n-2)+\Delta c_{1}\cdot n(n-2)/4=\Lambda.$
where in the last inequality we used Proposition˜4.7. ∎
@@ 5 Lower Bounds in Higher Dimension
In this section, we prove conditional lower bounds for approximating $\mathrm{EMDuT}$ with the $L_{1}$ or $L_{\infty}$ norm. Our lower bounds assume the popular Exponential Time Hypothesis (ETH), which postulates that the 3-SAT problem on $N$ variables cannot be solved in time $2^{o(N)}$ [ImpagliazzoP01].
@@
[…]
@@ 5.1 Lower Bound for $L_{1}$ Asymmetric
$\mathrm{EMD}_{1}(B_{i,j}+\tau,R_{i,j})+\mathrm{EMD}_{1}(B^{\prime}_{i,j}+\tau,R^{\prime}_{i,j})$ $\geq 2\min_{(u,v)\in E}\Big(|\tau_{i}-u|+|\tau_{j}-v|\Big)+(d-2)N$
$\geq(d-2)N+1.\qed$
We apply the Gadget Combination Lemma to the gadgets $B_{i,j},R_{i,j},B^{\prime}_{i,j},R^{\prime}_{i,j}$ for $1\leq i<j\leq d$. The value of the $\mathrm{EMDuT}_{1}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the above lemma implies the following. If $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k})\in[N]^{d}$ has a total cost of ${d\choose 2}\cdot(d-2)N\eqcolon\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Thus, each pair of gadgets contributes cost at least (d2)N(d-2)N, and at least one pair of gadgets contributes cost at least (d2)N+1(d-2)N+1, so the total cost is at least ${d\choose 2}\cdot(d-2)N+1=\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we had a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2})$, $1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N)$, and $d=k$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{3})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.
@@ 5.2 Lower Bound for $L_{1}$ Symmetric
[…]
We construct the following $2{k\choose 2}$ gadgets. For any $1\leq i<j\leq k$ we construct [^3] Here, we treat $B_{i,j}$ as a multi-set, containing |E|1|E|-1 times the points $q_{i,j}$. This can be avoided by adding a tiny perturbation to each copy, which makes $B_{i,j}$ a set without significantly changing any distances.
@@
Note that in any dimension $\ell\not\in\{i,j,i+k,j+k\}$ all points in $B_{i,j}$ and $B^{\prime}_{i,j}$ have coordinate 0, and thus the contribution of these dimensions to the total cost does not depend on the matching. For $(B_{i,j},R_{i,j})$ such a dimension $\ell$ contributes $|E|\cdot|\tau_{\ell}|\geq|E|\tau_{\ell}$ to the total cost, and for $(B^{\prime}_{i,j},R^{\prime}_{i,j})$ it contributes $|E|\cdot|N-\tau_{\ell}|\geq|E|(N-\tau_{\ell})$. Summing up over both gadgets and all $\ell\not\in\{i,j,i+k,j+k\}$ yields total cost at least (d4)|E|N(d-4)|E|N, with equality if $\tau\in[N]^{d}$.
@@
where we used $|x|\geq x$. Furthermore, equality holds if $\tau\in[N]^{d}$. Analogously, one can show that projected onto the dimensions $i,j,i+k,j+k$ the $L_{1}$ distance from $q_{i,j}+\tau$ to $\bar{p}_{i,u,j,v,N}$ is
@@
By summing up everything, and noting that the terms $\tau_{i}+\tau_{j}-\tau_{i+k}-\tau_{j+k}$ and $\tau_{i}-\tau_{j}+\tau_{i+k}+\tau_{j+k}$ cancel, we obtain
@@
We apply the Gadget Combination Lemma to the gadgets $B_{i,j},R_{i,j},B^{\prime}_{i,j},R^{\prime}_{i,j}$ for $1\leq i<j\leq k$. The value of the $\mathrm{EMDuT}_{1}$ of the resulting point sets $B,R$ is the sum of the costs of the gadgets. Hence, the above lemma implies the following. If $G$ has a $k$-Clique $v_{1},\ldots,v_{k}$, then $\tau\coloneqq(v_{1},\ldots,v_{k},v_{1},\ldots,v_{k})\in[N]^{d}$ has a total cost of ${k\choose 2}\cdot((d+4)|E|-8)N=:\Lambda$. On the other hand, if $G$ has no $k$-Clique, then for any $\tau\in\mathbb{R}^{d}$ there exist $1\leq i<j\leq k$ with $(\lfloor\tau_{i}\rceil,\lfloor\tau_{j}\rceil)\not\in E$ (as otherwise $(\lfloor\tau_{1}\rceil,\ldots,\lfloor\tau_{k}\rceil)$ would form a $k$-Clique). Thus, each pair of gadgets contributes cost at least ((d+4)|E|8)N((d+4)|E|-8)N, and at least one pair of gadgets contributes cost at least ((d+4)|E|8)N+1((d+4)|E|-8)N+1, so the total cost is at least ${k\choose 2}\cdot((d+4)|E|-8)N+1=\Lambda+1$.
For any $\varepsilon<1/\Lambda$, a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ could distinguish cost at most $\Lambda$ and cost at least $\Lambda+1$, and thus would solve the $k$-Clique problem. Hence, if we had a $(1+\varepsilon)$-approximation algorithm for $\mathrm{EMDuT}_{1}$ running in time $(n/\varepsilon)^{o(d)}$, then by setting $\varepsilon\coloneqq 0.9/\Lambda$ and observing $n=\mathcal{O}(N^{2}),1/\varepsilon=\mathcal{O}(\Lambda)=\mathcal{O}(N^{3})$, and $d=\mathcal{O}(k)$, we would obtain an algorithm for $k$-Clique running in time $(n/\varepsilon)^{o(d)}=\mathcal{O}(N^{5})^{o(k)}=N^{o(k)}$, which contradicts ETH by Theorem 5.2.